Problem 5-41A (Part Level Submission)

Hy and van Lamsweerde is a public accounting firm that offers two primary services, auditing and tax return preparation. A controversy has developed between the partners of the two service lines regarding who is contributing the greater amount to the bottom line. The area of contention is the assignment of overhead. The tax partners argue for assigning overhead on the basis of 40% of direct labour dollars, while the audit partners argue for implementing activity-based costing. The partners agree to use next year’s budgeted data for analysis and comparison. The following overhead data are collected:

Expected Use of Cost
Drivers per Service
Activity Cost Pools Cost Drivers Estimated
Overhead
Expected Use
of Cost Drivers
Audit Tax
Employee training Direct labour dollars $224,730 $1,872,753 $996,000 $876,753
Typing and secretarial Number of reports and forms 82,795 2,900 700 2,200
Computing Number of minutes 212,936 61,900 26,000 35,900
Facility rental Number of employees 145,740 42 22 20
Travel Per expense reports 82,900 Direct 56,700 26,200
749,101

1. Using activity-based costing, prepare a schedule showing the calculations of the activity-based overhead rates (per cost driver). (Round answers to 2 decimal places, e.g. 15.25.)

Activity-based overhead rates
Employee training $

per DL dollar
Secretarial $

per report/form
Computing $

per minute
Facility rental $

per employee
Travel $



2. Prepare a schedule assigning each activity’s overhead cost pool to each service based on the use of the cost drivers. (Round answers to 0 decimal places, e.g. 1525.)

Audit Tax
Training $

$

Secretarial

Computing

Rent

Travel (Direct)

Total overhead assigned $

$
Answer #1

1. Activity-based overhead rates = Estimated overhead/Expected use of cost drivers

Activity-based overhead rates
Employee training 0.12 per DL dollar
Secretarial 28.55 per report/form
Computing 3.44 per minute
Facility rental 3470.00 per employee
Travel

Working:

Activity-based overhead rates
Employee training =224730/1872753 per DL dollar
Secretarial =82795/2900 per report/form
Computing =212936/61900 per minute
Facility rental =145740/42 per employee

2.

Audit Tax
Training 119520 105210
Secretarial 19985 62810
Computing 89440 123496
Rent 76340 69400
Travel (Direct) 56700 26200
Total overhead assigned 361985 387116

Working:

Audit Tax
Training =996000*0.12 =876753*0.12
Secretarial =700*28.55 =2200*28.55
Computing =26000*3.44 =35900*3.44
Rent =22*3470 =20*3470
Travel (Direct) 56700 26200